in how many ways can 6 people apply for 2 different jobs?
Answer by Edwin McCravy(20054)

in how many ways can 6 people apply for 2 different jobs?
The answer would depend of whether they can apply for both jobs 
or whether they can only apply for 1 of the jobs.

I'll do it both ways.

Assuming they can do any of 3 things --
apply for either J only,  K only, or both J and K

There are 3 things A can do, apply for J, K, or J&K

For each of the 3 things A can do, there are 3 things B can do.
That's 3×3=32=9 things A and B can do. 

For each of the 3×3=32=9 things A and B can do, there 
are 3 things C can do.

That's 3×3×3=33=27 things A,B, and C can do. 

... etc. ...

For each of the 3×3×3×3×3=35=243 things A,B,C,D and E can do, 
there are 3 things F can do.

That's 3×3×3×3×3×3=36=729 things A,B,C,D,E, and F can do.  

Answer: 36=729

-----------------

Assuming they can only do either of 2 things --
apply for either J only or apply for K only. 
[They can't apply for both jobs.]

There are 2 things A can do, apply to J or K

For each of the 2 things A can do, there are 2 things B can do.
That's 2×2=32=4 things A and B can do. 

For each of the 2×2=22=4 things A and B can do, there 
are 2 things C can do.

That's 2×2×2=23=8 things A,B, and C can do. 

... etc. ...

For each of the 2×2×2×2×2=25=32 things A,B,C,D and E can do, 
there are 2 things F can do.

That's 2×2×2×2×2×2=26=64 things A,B,C,D,E, and F can do.  

Answer: 26=64

---------------

Edwin